# The Telephone Game: Evaluating Semantic Drift in Unified Models

### v1

@@ Abstract
Employing a single, unified model (UM) for both visual understanding (image-totext: I2T) and and visual generation (text-to-image: T2I) has opened a new direction in Visual Language Model (VLM) research. While UMs can also support broader unimodal tasks (e.g., text-to-text, image-to-image), we focus on the core cross-modal pair T2I and I2T, as consistency between understanding and generation is critical for downstream use. Existing evaluations consider these capabilities in isolation: FID and GenEval for T2I, and benchmarks such as MME, MMBench for I2T. These single-pass metrics do not reveal whether a model that “understands” a concept can also “render” it, nor whether meaning is preserved when cycling between image and text modalities. To address this, we introduce the Unified Consistency Framework for Unified Models (UCF-UM), a cyclic evaluation protocol that alternates I2T and T2I over multiple generations to quantify semantic drift. UCF formulates 3 metrics: (i) Mean Cumulative Drift (MCD), an embedding-based measure of overall semantic loss; (ii) Semantic Drift Rate (SDR), that summarizes semantic decay rate; and (iii) Multi-Generation GenEval (MGG), an object-level compliance score extending GenEval. To assess generalization beyond COCO, which is widely used in training; we create a new benchmark ND400, sampled from NoCaps and DOCCI and evaluate on seven recent models. UCF-UM reveals substantial variation in cross-modal stability: some models like BAGEL maintain semantics over many alternations, whereas others like Vila-u drift quickly despite strong single-pass scores. Our results highlight cyclic consistency as a necessary complement to standard I2T and T2I evaluations, and provide practical metrics to consistently assess unified model’s cross-modal stability and strength of their shared representations. Code will be provided in this repository: https://github.com/mollahsabbir/Semantic-Drift-in-Unified-Models
@@ 1 Introduction
Figure 1: (a) Illustrates image generation and image understanding functionalities of a unified model. (b) Example evaluation using proposed UCF-UM framework. In this example the unified model starts from a textual prompt $T^{(0)}$ about a suitcase and a banana. After successive T2I and I2T steps we observe in generation 5 the model fails to generate a convincing suitcase, and subsequently the suitcase disappeared from future generations. Also, at generation 5 it created two bananas instead of one, which culminated in lots of bananas in generation 7 and forward.
@@
Multimodal Unified Models (UMs) combine visual understanding and generation within a single framework, enabling them to perform a wide range of unimodal tasks (e.g., text-to-text, image-to-image) as well as cross-modal tasks (e.g., image-to-text, text-to-image). By sharing representations across modalities, UMs can demonstrate interesting emerging capabilities such as intelligent photo editing, e.g. BAGEL [8]. However, despite rapid model progress, UM evaluation remains fragmented. Existing metrics assess image understanding and image generation in isolation; e.g. MME, MMBench, POPE, VQA [9, 14, 12, 1] are used for evaluating understanding (I2T), and Inception score, CLIPScore, FID, GenEval [20, 11, 22] are used for evaluating image synthesis (T2I), while overlooking the retention of important information during T2I or I2T multi-turn conversion. In other words, current single-pass metrics do not assess the retention of entities, attributes, relations, and counts under alternating I2T $\leftrightarrow$ T2I conversions. We defer unimodal tasks and center our analysis on I2T and T2I tasks as the potential for semantic divergence and its impact on real use is most pronounced on the cross-modal tasks.
Fig. 1 demonstrates this limitation. Much like the popular children’s game called Telephone Game, where a whispered message drifts in meaning as it passes from person to person, UMs tend to lose or distort semantic meaning when cycling between text and image representations as shown in Fig. 1(b). Starting from a textual prompt: “a suitcase left of a banana”, the model produces an image $I^{(1)}$ correctly, which is then captioned (I2T) to form the next prompt $T^{(2)}$, and so on. Although each individual step can look plausible in isolation, semantic drift accumulates across the cycles: by generation 5, the suitcase is no longer convincingly rendered and soon disappears entirely; meanwhile, the banana count inflates from one, to two, and eventually to many. Notably, a model may score well on single-pass I2T or T2I metrics, while still exhibiting these cross-modal inconsistencies, which the current metrics fail to capture.
@@
Figure 2: An example of unified inconsistency in the BAGEL unified model. Given an image of a chess board along with a question (top), BAGEL performs I2T correctly answering “white side wins”. By creating another caption for the T2I prompt (bottom), BAGEL should generate a chess board image consistent with the same semantic predicate (white winning side). However, the model generates a generic, mismatched chessboard image. This exposes a UM inconsistency: BAGEL’s correct visual reasoning (I2T) does not carry over to generation (T2I) for the concept “winning side in chess”.
@@
There are several ways to evaluate a model’s image generation capabilities. For example, ClipScore [10] uses clip embeddings to measure semantic alignment of the prompt with generated images. However, it strongly relies on clip embeddings, which may not always be reflective with human perceptions [22]. Fréchet Inception Distance (FID) [11] measures the distributional similarity between the generated images and real images, but ignores the generated image’s faithfulness to the input prompt. A model that ignores the input text and produces high-quality yet off-prompt images can still score well [22]. GenEval [22] improves on prompt alignment by checking object and relation-level compliance with detection models, but, by design, does not assess overall visual quality or realism, and like FID, remains a single-pass measure.
A similar limitation appears on the image-understanding benchmarks such as MME and MMBench [9, 14] which assess I2T skills in isolation, without testing whether the model’s understanding capability aligns with its generation capability. As illustrated in Fig. 2, even state-of-the-art unified models like BAGEL [8] can correctly reason about a chessboard image in I2T identifying that “the white side wins” yet fail to produce a faithful T2I image of the same winning scenario. This mismatch highlights the need for metrics that directly measure semantic drift across modalities rather than a single pass metric.
@@
To address this gap, we propose the Unified Consistency Framework for Unified Models (UCF-UM), a cyclic evaluation protocol designed to quantify how well UMs preserve semantic meaning under repeated T2I and I2T conversions. Starting from an initial input $T^{(0)}$ (text) or $I^{(0)}$ (image), the model alternates T2I or I2T to produce a sequence $\{I^{(g)},T^{(g)}\}$, where g denotes generation step. At each generation g, UCF-UM measures semantic similarity back to the initial input and across steps, capturing drift directions and exposing misalignment between a model’s understanding and generation spaces. We employ CLIP [19], DINO [4], and MPNet [26] embeddings for text–image, image–image, and text–text comparisons, respectively. For rigorous testing, we design three different metrics: Mean Cumulative Drift (MCD), Semantic Drift Rate (SDR), and Multi-Generation Geneval (MGG). In MCD, we use raw embedding distance scores to quantify cumulative information retention, SDR tells us properties of the drift through parameters such as decay rate, and MGG extends the GenEval benchmark for multiple generations. We propose a benchmark dataset ND400, sampling 200 image-text pairs from NoCaps [2] and 200 image-text pairs DOCCI [17] datasets. These two datasets were selected for their novel objects and fine-grained visual details that better probe generalization. We benchmark $7$ recent models spanning shared-weight, partially shared, and decoupled architectures, to analyze how architectural design choices influence semantic stability.
@@
Our experiments reveal substantial variation in semantic drift behavior across models. For example, BAGEL [8] maintains strong semantic fidelity across multiple generation cycles, whereas models like Vila-U [30] and Janus [29] degrade rapidly, exposing weaker coupling between their visual understanding and visual generation capabilities despite competitive single-pass metrics. These findings underscore the need to move beyond isolated I2T or T2I metrics and toward evaluations that directly measure cross-modal consistency.
@@
  • We formalize semantic drift problem and show that single-pass understanding or generation metrics cannot expose gaps between a model’s visual understanding and its image generation capabilities.
  • We propose the Unified Consistency Framework for Unified Models (UCF-UM), which jointly evaluates image understanding (I2T) and image generation (T2I) through 3 different metrics, and tracks semantic preservation over multiple cross-modal transitions.
  • We further extend GenEval [22] to a multi-generation setting, enabling it to capture semantic preservation beyond single prompt–generation pairs. In this setup, overall performance differences between models amplifies and becomes clearly observable.
@@ 2 Unified Models
T2I generation has advanced with diffusion-based models such as DALL·E 2 [21], Imagen [24], and Stable Diffusion [23], which synthesize high-fidelity images from textual prompts. Image captioning, on the other hand, has evolved from CNN-RNN pipelines [25] to transformer-based decoders [7, 13] trained with large web-scale data. Recent works in unified models have started investigating how to unite understanding and generation under one architecture.
Unified models employ visual and textual modalities as both input and output. The motivation is that these universal models facilitate richer semantic interoperability among the two tasks, I2T and T2I. Also, having a single architecture can reduce training time, and simplify deployment. Chameleon [27] is one of the early works in this domain which aimed to auto-regressively generate text tokens and image embeddings. Later, Transfusion [34] fused the auto-regressive and diffusion loss within a single architecture. Show-o [31] has also used two different objectives, next token prediction for text generation, and masked token prediction [5] for image generation. Vila-u [30] uses next token prediction with different text and vision decoders. Janus and Janus-pro [29] employ separate encoders for image input during understanding and generation. The idea is that a model might require different level of information for understanding and generation. Other works like Blip-3o [6] demonstrates good quality of image generation by leveraging a separate diffusion transformer head. A recent work, BAGEL [8] demonstrates some unique capabilities of unified models by training on a large-scale interleaved dataset.
@@
This category has received the most attention in recent research. These models leverage a single model, typically a transformer decoder, to perform a wide spectrum of unimodal and cross-modal tasks, with T2I and I2T generation being prominent examples. The encoder component can vary where some models employ a shared visual encoder across tasks, while others use distinct encoders for generation and understanding. In our experiments, we use $5$ such models: BAGEL [8], Janus 1.3B [29], Janus Pro 7B [29], Show-o [31], and Vila-u [30].
@@
Models in this category retain a degree of parameter sharing, while delegating specific responsibilities to task-specific modules. This design allows more flexibility in handling modality-specific complexities while preserving shared knowledge across tasks. We use Blip-3o [6] which incorporates a dedicated diffusion model for image generation.
@@
Models in the third category are formed by constructing a unified pipeline by composing independently trained models, which in tandem can emulate unified behavior. The example we have used is pairing a VLM like LLaVA[13] for I2T with a Stable Diffusion [18] model for T2I. This setup enables task interoperability without requiring joint training or weight sharing.
Figure 4: Unified Consistency Evaluation framework (UCF-UM). We alternate between text-to-image (T2I) and image-to-text (I2T) generations in two setups: Text-First-Chain (a) and Image-First-Chain (b). Blue arrows denote I2T; purple arrows denote T2I; dashed black arrows indicate similarities computed back to the initial input in both same‑ and cross‑modality directions used for MCD. Across generations, concepts drift despite plausible single steps: a “red F-450 truck” evolves into a semi‑truck with changing attachments and positions; in the image‑first chain, group size inflates and new objects (e.g. a sports ball) appear. The proposed cyclic evaluation reveals cross‑modal concept drift that single‑pass metrics overlook, enabling direct comparison of unified model’s semantic stability.
3 Unified Consistency Evaluation
We propose a cyclic evaluation framework UCF-UM which provides three different metrics to measures how well a unified model preserves semantic fidelity when alternating between I2T and T2I. As illustrated by the single alternation in Fig. 2 where correct I2T reasoning about a chessboard does not translate into a faithful T2I rendering; UCF-UM generalizes this intuition to multi-generation cycles and provides quantitative measures of semantic drift. This framework considers how quickly the information from the initial input is drifted in subsequent generations. In this setting, we treat the $\mathcal{UM}$ as a model composed of at least two functionalities.
  • Image Generation (T2I): $\mathcal{UM}_{\mathrm{\texttt{T2I}}}:\mathcal{T}\rightarrow\mathcal{I}$, which synthesizes an image given a textual description.
  • Image Understanding (I2T): $\mathcal{UM}_{\mathrm{\texttt{I2T}}}:\mathcal{I}\rightarrow\mathcal{T}$, which generates a textual description from a given image and a prompt (e.g.’Describe this image’)
Here, $\mathcal{T}$ denotes the set of all possible text representations (e.g., captions, instructions), and $\mathcal{I}$ denotes the set of all possible image representations. Let $\mathcal{D}=\{(I_{i},T_{i})\}_{i=1}^{N}$ represent a dataset of $N$ paired samples, where each $I_{i}\in\mathcal{I}$ and each $T_{i}\in\mathcal{T}$ is its corresponding caption. A generation step is defined as the application of either $\mathcal{UM}_{\mathrm{\texttt{T2I}}}$ or $\mathcal{UM}_{\mathrm{\texttt{I2T}}}$ to transform an input from one modality into the other. We define alternating chains of length G starting from either text or image. Let $g\in\{0,1,\ldots,G\}$ be the generation step index. Then similar to the chains defined in [3], we consider two experimental setups depending on the initial modality:
@@ 3 Unified Consistency Evaluation
The intuition for UCF-UM is that a semantically consistent model will preserve the core meaning of the original content across many generations of alternating T2I and I2T; A weaker model will drift away from the original meaning more quickly. To systematically measure this degradation, in our framework we propose three distinct metrics. MCD provides a holistic measure of drift based on embedding similarity. SDR offers a granular analysis of the decay rate by fitting it to a power-law function. Finally, MGG grounds the evaluation in object-level fidelity by extending the GenEval benchmark across multiple generations.

### v2

@@ Abstract
Employing a single, unified model (UM) for both visual understanding (image-to-text: I2T) and visual generation (text-to-image: T2I) has opened a new direction in Visual Language Model (VLM) research. While UMs can also support broader unimodal tasks (e.g., text-to-text, image-to-image), we focus on the core cross-modal pair T2I and I2T. Existing evaluation benchmarks consider these capabilities in isolation: FID and GenEval for T2I, and benchmarks such as MME, MMBench for I2T. These isolated single-pass metrics do not reveal cross-consistency: whether a model that “understands” a concept can also “render” it, nor whether semantic meaning is preserved when cycling between image and text modalities. To address this, we introduce the Semantic Drift Protocol (SDP) for Unified Models, a cyclic evaluation protocol that alternates I2T and T2I over multiple generations to quantify semantic drift. We propose two metrics: (i) Mean Cumulative Drift (MCD), an embedding-based measure of overall semantic drift; and (ii) Multi-Generation GenEval (MGG), an object-level compliance score extending GenEval. To assess generalization beyond COCO dataset, which is widely used in training; we create a new benchmark Nocaps+Docci400, sampled from NoCaps and DOCCI and evaluated on seven recent models. SDP reveals substantial variation in cross-modal stability: some models like BAGEL maintain semantic meaning over many alternations, whereas others like VILA-U drift quickly despite strong single-pass scores. Our results highlight SDP as a necessary complement to standard I2T and T2I evaluations. Code is available at https://github.com/mollahsabbir/Semantic-Drift-in-Unified-Models
Figure 1: (a) Illustrates image generation and image understanding functionalities of a unified model. (b) Telephone Game: We propose a new form of evaluation consisting of alternating T2I and I2T steps. Here, the unified model starts from a textual prompt $T^{(0)}$ about a suitcase and a banana. At every step we observe semantic drift. For example, in the 5th generation, the model fails to generate a convincing suitcase, which also hints at cross-inconsistency. These phenomena are magnified under the multi-generation telephone game evaluation, allowing it to capture more subtle performance differences between models.
@@ 1 Introduction
Multimodal Unified Models (UMs) combine visual understanding and generation within a single framework, enabling a wide range of unimodal tasks (e.g., text-to-text, image-to-image) as well as cross-modal tasks (e.g., image-to-text, text-to-image). By sharing representations across modalities, UMs can demonstrate interesting emerging capabilities such as intelligent photo editing, e.g. BAGEL Deng et al. (2025). Despite rapid model progress, UM evaluation remains fragmented. Existing metrics assess image understanding and image generation in isolation; e.g., MMEFu et al. (2024), MMBench Liu et al. (2024), POPE Li et al. (2023b), VQA Agrawal et al. (2016) are used for evaluating understanding (I2T), and Inception score Radford et al. (2016), CLIPScore Hessel et al. (2022), FID Heusel et al. (2017), GenEval Ghosh et al. (2023) are used for evaluating image synthesis (T2I), while overlooking the retention of important information during T2I or I2T multi-turn conversion. In other words, current single-pass metrics do not assess the retention of entities, attributes, relations, and counts under alternating I2T $\leftrightarrow$ T2I conversions. We defer unimodal tasks and center our analysis on I2T and T2I tasks as the potential for semantic divergence and its impact on real use is most pronounced on the cross-modal tasks.
@@
Figure 2: An example of cross-consistency in the BAGEL unified model. Given an image of a chess board along with a question (top), BAGEL performs I2T, correctly answering “white side wins”. By creating another caption for the T2I prompt (bottom), BAGEL should generate a chess board image consistent with the same semantic predicate (white winning side). However, the model generates a generic, mismatched chessboard image. This exposes a unified model inconsistency: BAGEL’s correct visual reasoning (I2T) does not carry over to generation (T2I) for the concept “winning side in chess”.
We begin by formalizing two key notions, “semantic-drift” and “ cross-consistency”. Semantic drift is the loss or distortion of meaning that accumulates when an input is repeatedly transformed across modalities via T2I and I2T. Cross-consistency refers to the overlap between what a model can generate as images from text and what it can faithfully understand from images as text. Much like the popular children’s game called Telephone Game, where a whispered message drifts in meaning as it passes from person to person, UMs tend to lose or distort semantic meaning when cycling between text and image representations as shown in Fig. 1(b). Starting from a textual prompt: “a suitcase left of a banana”, the model produces an image $I^{(1)}$ correctly, which is then captioned (I2T) to form the next prompt $T^{(2)}$, and so on. Although each individual step can look plausible in isolation, semantic drift accumulates across the cycles: by generation 5, the image has changed drastically. Notably, a model may score well on isolated single-pass I2T or T2I metrics, while still exhibiting these cross-modal inconsistencies, which the current metrics fail to capture. The concept of cross-consistency is illustrated in Fig. 2, where even state-of-the-art unified models like BAGEL Deng et al. (2025) can correctly reason about a chessboard image in I2T identifying that “the white side wins”, yet fail to produce a faithful T2I image of the same winning scenario.
@@
There are several ways to evaluate a model’s image generation capabilities. For example, ClipScore Hessel et al. (2022) uses clip embeddings to measure semantic alignment of the prompt with generated images. However, it strongly relies on clip embeddings, which may not always be reflective with human perceptions Ghosh et al. (2023). Fréchet Inception Distance (FID) Heusel et al. (2017) measures the distributional similarity between the generated images and real images, but ignores the generated image’s faithfulness to the input prompt. A model that ignores the input text and produces high-quality, yet off-prompt images can still score well Ghosh et al. (2023). GenEval Ghosh et al. (2023) improves on prompt alignment by checking object and relation-level compliance with detection models, however, by design, does not assess overall visual quality or realism, and like FID, remains a single-pass measure. A similar limitation is observed in the image-understanding benchmarks, such as MME Fu et al. (2024) and MMBench Liu et al. (2024) which assess I2T skills in isolation, without testing whether the model’s understanding capability aligns with its generation capability.
@@
To address this gap, we evaluate unified models in single and multi-pass settings respectively. In the single-pass setting (one-step I2T and T2I on paired image–caption data), we perform a human cross-consistency study to judge consistency between model outputs relative to its inputs. We ask human annotators to rank the model outputs, and also categorize each output in one of good, medium, or poor fidelity. The fidelity scores indicate the degree to which inconsistencies are present, while the rankings establish relative model performance according to human judgment. These results show the cross-consistency in the evaluated models. In multi-pass evaluation, we propose the Semantic Drift Protocol for Unified Models (SDP), a cyclic evaluation protocol designed to quantify how well UMs preserve semantic meaning under repeated T2I and I2T conversions. Starting from an initial input $T^{(0)}$ (text) or $I^{(0)}$ (image), the model alternates T2I or I2T to produce a sequence $\{I^{(g)},T^{(g)}\}$, where $g$ denotes generation step. At each generation $g$, SDP measures semantic similarity back to the initial input and across steps, capturing drift directions and exposing misalignment between a model’s understanding and generation spaces. We employ CLIP Radford et al. (2021), DINO Caron et al. (2021), and MPNet Song et al. (2020) embeddings for text–image, image–image, and text–text comparisons, respectively. For rigorous testing, we design two different metrics: Mean Cumulative Drift (MCD), and Multi-Generation Geneval (MGG). In MCD, we use raw embedding distance scores to quantify cumulative information retention, and MGG extends the GenEval benchmark for multiple generations. We propose a new benchmark dataset Nocaps+Docci400, sampling 200 image-text pairs from NoCaps Agrawal et al. (2019) and 200 image-text pairs DOCCI Onoe et al. (2024) datasets. These two datasets were selected for their novel objects and fine-grained visual details that better probe generalization. We benchmark $7$ recent models spanning shared-weight, partially shared, and decoupled architectures, to analyze how architectural design choices influence semantic stability.
@@
Our experiments reveal substantial variation in semantic drift behavior across models. For example, BAGEL Deng et al. (2025) maintains strong semantic fidelity across multiple generation cycles, whereas models like VILA-U Wu et al. (2025) and Janus Wu et al. (2024) degrade rapidly, exposing weaker coupling between their visual understanding and visual generation capabilities despite competitive single-pass metrics. These findings underscore the need to move beyond isolated I2T or T2I metrics and toward evaluations that directly measure cross-consistency.
@@
  • We formalize the cross-consistency and semantic drift problem, showing that single-pass metrics cannot expose gaps between a model’s understanding and generation capabilities.
  • We propose the Semantic Drift Protocol (SDP), which jointly evaluates I2T and T2I over multiple transitions to track semantic preservation.
  • We extend GenEval Ghosh et al. (2023) to a multi-generation setting, which amplifies observable performance differences between models.
  • We conduct a human study to determine cross-consistency in existing models and provide a comparative ranking.
@@ 2 Unified Models
Unified models employ visual and textual modalities as both input and output. The motivation is that these universal models facilitate richer semantic interoperability among the two tasks, I2T and T2I.
@@
This category has received the most attention in recent research. These models leverage a single model, typically a transformer decoder, to perform a wide spectrum of unimodal and cross-modal tasks, with T2I and I2T generation being prominent examples. The encoder component can vary where some models employ a shared visual encoder across tasks, while others use distinct encoders for generation and understanding. In our experiments, we use $5$ such models: BAGEL Deng et al. (2025), Janus 1.3B Wu et al. (2024), Janus Pro 7B Wu et al. (2024), Show-o Xie et al. (2024), and VILA-U Wu et al. (2025).
@@
Models in this category retain a degree of parameter sharing, while delegating specific responsibilities to task-specific modules. This design allows more flexibility in handling modality-specific complexities while preserving shared knowledge across tasks. We use BLIP3-o Chen et al. (2025) which incorporates a dedicated diffusion model for image generation.
@@
Models in the third category are formed by composing independently trained components for I2T and T2I, yielding a unified pipeline that emulates end-to-end behavior. In this setup, the I2T component is handled by pretrained Multimodal LLMs ( Li et al. (2023a); Sirnam et al. (2024); Liu et al. (2023)), while the T2I direction is addressed by generative diffusion models ( Rombach et al. (2022); Podell et al. (2023); Ramesh et al. (2022); Chen et al. (2024)). For our experiments, we pair LLaVA Liu et al. (2023) for I2T with a SDXL Podell et al. (2023) model for T2I. This setup enables task interoperability without requiring joint training or weight sharing.
3 Semantic Drift Evaluation
We propose a cyclic evaluation Protocol SDP which provides two different metrics to measures how well a unified model preserves semantic fidelity when alternating between I2T and T2I. SDP proposes to evaluate on multi-generation cycles to provide quantitative measures of semantic drift. In this setting, we treat the $\mathcal{UM}$ as a model composed of at least two functionalities. Image Generation: $\mathcal{UM}_{\mathrm{\texttt{T2I}}}:\mathcal{T}\rightarrow\mathcal{I}$, which synthesizes an image given a textual description. Image Understanding (I2T): $\mathcal{UM}_{\mathrm{\texttt{I2T}}}:\mathcal{I}\rightarrow\mathcal{T}$, which generates a textual description from a given image. Here, $\mathcal{T}$ denotes the set of all possible text representations (e.g., captions, instructions), and $\mathcal{I}$ denotes the set of all possible image representations.
Let $\mathcal{D}=\{(I_{i},T_{i})\}_{i=1}^{N}$ represent a dataset of $N$ paired samples, where each $I_{i}\in\mathcal{I}$ and each $T_{i}\in\mathcal{T}$ is its corresponding caption. A generation step is defined as the application of either $\mathcal{UM}_{\mathrm{\texttt{T2I}}}$ or $\mathcal{UM}_{\mathrm{\texttt{I2T}}}$ to transform an input from one modality into the other. We define alternating chains of length $G$ starting from either text or image. Let $g\in\{0,1,\ldots,G\}$ be the generation step index. Then similar to the chains defined in Bahng et al. (2025), we consider two experimental setups depending on the initial modality:
@@ 3 Semantic Drift Evaluation
Figure 4: Semantic Drift Protocol (SDP). We alternate between text-to-image (T2I) and image-to-text (I2T) generations in two setups: Text-First-Chain (a) and Image-First-Chain (b). Blue arrows denote I2T; purple arrows denote T2I; dashed black arrows indicate similarities computed back to the initial input in both same‑ and cross‑modality directions used for MCD. Across generations, concepts drift despite plausible single steps: a “red F-450 truck” evolves into a semi‑truck with changing attachments and positions; in the image‑first chain, group size inflates and new objects (e.g., a sports ball) appear. The proposed cyclic evaluation reveals cross‑modal concept drift that single‑pass metrics overlook, enabling direct comparison of unified model’s semantic stability.
@@
The intuition for SDP is that a semantically consistent model will preserve the core meaning of the original content across many generations of alternating T2I and I2T; A weaker model will drift away from the original meaning more quickly. To systematically measure this degradation, in our protocol we propose two distinct metrics. MCD provides a holistic measure of drift based on embedding similarity. On the other hand, MGG grounds the evaluation in object-level fidelity by extending the GenEval benchmark across multiple generations.